# Separate Source Channel Coding Is Still What You Need: An LLM-based Rethinking

### v1

@@ Abstract
Along with the proliferating research interest in Semantic Communication (SemCom), Joint Source Channel Coding (JSCC) has dominated the attention due to the widely assumed existence in efficiently delivering information semantics. Nevertheless, this paper challenges the conventional JSCC paradigm, and advocates for adoption of Separate Source Channel Coding (SSCC) to enjoy the underlying more degree of freedom for optimization. We demonstrate that SSCC, after leveraging the strengths of Large Language Model (LLM) for source coding and Error Correction Code Transformer (ECCT) complemented for channel decoding, offers superior performance over JSCC. Our proposed framework also effectively highlights the compatibility challenges between SemCom approaches and digital communication systems, particularly concerning the resource costs associated with the transmission of high precision floating point numbers. Through comprehensive evaluations, we establish that empowered by LLM-based compression and ECCT-enhanced error correction, SSCC remains a viable and effective solution for modern communication systems. In other words, separate source and channel coding is still what we need!
@@ I Introduction
Semantic communication (SemCom) has garnered significant attention in recent years, with researchers exploring innovative approaches to enhance the efficiency and reliability of information transmission [1]. Generally, SemCom leverages deep learning-based Joint Source-Channel Coding (JSCC) methods to preserve global semantic information and local texture during the transmission process. DeepJSCC [2] pioneers these works by implementing JSCC with feedback and allowing for real-time adaptation to channel conditions. Along with its steady progress, JSCC has been substantially studied, mostly with the optimization objective shifting from bit error rates to the semantic relevance of the transmitted information in SemCom [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. However, albeit the awfully exploded research interest, one critical question remains unsolved, that is, why does the joint approach stand out, as a separate source channel coding (SSCC) shall promise more degree of freedom from an optimization perspective?
@@
As the terminology implies, SSCC encompasses two decoupled ingredients, namely source coding and channel coding. The former part lies in effectively compressing the context, and the effectiveness of underlying Deep Neural Networks (DNN)-based predictors, such as Recurrent Neural Networks (RNN)-based DeepZip [15], Long Short Term Memory (LSTM)-based [16, 17] and hybrid DNN-based Dzip [18], toward achieving acceptable text compression, have been validated widely. More prominently, Transformer-based [19] and Large Language Model (LLM)-based compression springs up recently [20, 21, 22, 23, 24]. The latest research [25] unveils the equivalence between compression and prediction. In other words, in the general framework where statistical models predict symbols and encoders use predictive probabilities to perform compression, better predictive models lead directly to better compressors [25]. Hence, the astonishing capability of LLM implies to yield an unprecedented source codec. On the other hand, Error Correction Code (ECC) plays an indispensable role in the channel coding. Although some advanced algebraic block codes like Bose–Chaudhuri–Hocquenghem (BCH) codes [26], Low-Density Parity-Check (LDPC) codes [27] and Polar codes [28], can somewhat ensure the reliability of transmission, the efficient decoding of ECC is an unresolved difficulty. Recently, DNNs start to demonstrate their contribution in channel coding. For example, deep learning models are implemented to achieve Belief Propagation (BP) decoding [29, 30, 31], while a model-free Error Correction Code Transformer (ECCT) for algebraic block codes [32] contributes to the enhancement of the decoding reliability.
@@
In this paper, we introduce an SSCC framework based on LLM and ECCT, and demonstrate the performance superiority over JSCC. This performance improvement primarily arises after tackling the underlying incompatibility between conventional SemCom approaches [3, 7, 8, 9, 10, 11, 12, 13, 14] and digital communication architectures [33]. Particularly, those approaches simply assume the deliverability of encoded semantic feature vectors while neglect the energy cost to transmit high-precision floating point numbers [33]. Unfortunately, further quantization [5, 6] and digital modulation can compromise the widely assumed existence of performance superiority in JSCC. Meanwhile, in contrast to the blunt, direct utilization of the astonishing semantic interpretation capability [34, 35, 36], the deployment of LLMs focuses on the compression and encoding of text to squeeze the largely untapped redundancy. Therefore, our work is also significantly different from those integrating Generative AI (GAI) and SemCom [37, 38, 39, 40, 41, 42, 43]. Furthermore, the adoption of ECCT boosts the effectiveness of SSCC under some cases. In a nutshell, a comprehensive evaluation of LLM and ECCT-based SSCC shows that separate source channel coding is still what we need!
@@
The rest of this paper is organized as follows. Section II introduces the SSCC system model while its key components are enumerated in Section III. Numerical results are presented in Section IV to show the performance superiority of our SSCC system. Finally, Section V concludes this paper with discussions on future works. For convenience, we list the major notations of this paper in Table I.
@@
Notation Defination
$\bm{s}_{1:N_{s}},\bm{\hat{s}}$ The transmitted text sequence and the recovered text sequence after source decoding at the receiver side
$\bm{t}_{1:N_{t}}$ The token sequence obtained by tokenizing the text sequence
@@
$\mathcal{D},D_{i},\tau$ The dictionary of source coder, the $i$-th character in the dictionary and vocabulary of the dictionary
$\mathbb{I}_{k},l_{k},u_{k}$ The probability interval in step $k$ of source coding and its corresponding lower and upper bounds
$\bm{m},\bm{\hat{m}}$ Message encoded by the source coder and the received (and channel decoded) message
$\lambda$ The probability interval, determined by the codeword, in a decimal form
$N,K$ The codeword length and message length of error correction code $C_{e}(N,K)$
$\bm{G,H}$ The generator matrix and the parity check matrix
$\bm{x},\bm{x}_{s},\bm{\hat{x}}$ Codeword encoded by the channel coder, the corresponding sign after BPSK modulation and the soft approximation of transmitted codeword after channel decoding
$\mathcal{N}(\cdot,\cdot),\sigma_{n}$ The Gaussian distribution and the standard deviation of noise
@@
$\bm{z},\bm{\tilde{z}},\bm{\hat{z}}$ The additive Gaussian noise, as well as its corresponding multiplicative noise and the prediction result by ECCT
$\bm{y},\bm{y}_{b},\bm{{\tilde{y}}}$ The noisy codeword, its binary form and the result of pre-processing noisy codeword
$\texttt{syn}(\cdot)$ The syndrome of codes defined in ECCT
$f(\cdot)$ The decoding function of ECCT
$\bm{W}$ The learnable embedding matrix for high-dimensional mapping
$\bm{g}(\cdot)$ The code aware self-attention mask
$R_{c},R_{m}$ The code rate and the modulation rate
@@ II System Model
  • Source Encoding: The input text sequence denoted as $\bm{s}_{1:N_{s}}$ undergoes a source encoder, which converts characters into a compressed binary message $\bm{m}\in\{0,1\}^{K}$. During source encoding, arithmetic coding (AC) can be leveraged for effective compression here, while for LLM-based processing, an intermediate result (i.e., a sequence of tokens $\bm{t}_{1:N_{t}}$) can be obtained during the transformation from $\bm{s}$ to $\bm{m}$.
  • Channel Encoding and Modulation: The message $\bm{m}$ is then encoded via a LDPC code $C_{e}(N,K)$, which is selected for its excellent error-correction capabilities and compatibility with iterative decoding algorithms, also mentioned in [32]. The encoding process employs a generator matrix $\bm{G}$ to transform the message in $\bm{m}$ to a codeword $\bm{x}\in\{0,1\}^{N}$. The parity check matrix $\bm{H}$, which satisfies $\bm{G}\cdot\bm{H}^{T}=0$ and $\bm{H}\cdot\bm{x}=0$ is a key component of the LDPC decoding process. Afterwards, BPSK modulation maps the binary codeword $\bm{x}$ to a sequence of symbols $\bm{x}_{s}\in\{0,1\}^{N}$ suitable for transmission over the wireless channel. Notably, other error correction codes such as Polar codes [28] can be applied as well.
@@
  • Demodulation and Channel Decoding: BPSK demodulation recovers a binary codeword $\bm{x}\in\{0,1\}^{N}$ from $\bm{x}_{s}$. Subsequently, LDPC decoding via the bit-flipping algorithm reconstructs the original message $\bm{m}\in\{0,1\}^{K}$ from $\bm{x}$. Before channel decoding, some complementary procedure such as ECCT can be applied to enhance the decoding performance.
  • Source Decoding: The recovered message $\bm{\hat{m}}$ is ultimately decoded by the source decoder, which reconstructs the original text sequence $\bm{s}$ from the message, effectively reversing the encoding process. Similar to the encoder, the decoder can implement an arithmetic decoding.
In comparison, JSCC often leverages an end-to-end DNN to implement source and channel codecs. Here, the terminology “end-to-end” implies the generally joint training of source and channel codes in most works. The details of JSCC can be found in [1] and the references therein. In the next section, we will address how to leverage the strength of LLM for enhanced text compression and reconstruction, combined with the robustness of ECCT-complemented LDPC codes for error correction, which is shown in Fig. 1.
Figure 1: Framework of LLM-based and ECCT-complemented SSCC system.
@@ III-A LLM-based Source Coding
Given the source distribution $\rho$, lossless compression aims to encode a text sequence $\bm{s}_{1:N_{s}}$ sampled from $\rho$ into a binary code $\bm{m}=C_{s}(\bm{s}_{1:N_{s}})$ of minimal possible length $l_{C_{s}}(\bm{s}_{1:N_{s}})$ with no loss of original information. According to Shannon’s source coding theorem [44], the optimal expected bits is $L_{\min}=\mathbb{E}_{\bm{s}\sim\rho}[-\log_{2}\rho(\bm{s})]$. To obtain a code with such optimal length, arithmetic coding [45, 46], a form of entropy encoding, is typically adopted contingent on a probabilistic model over $\rho$ or its marginal distribution. Arithmetic coding implies that frequently used characters be stored with fewer bits while rarely occurred characters correspond to more bits, resulting in fewer bits used in total.
In particular, the input text sequence denoted as $\bm{s}_{1:N_{s}}$ undergoes tokenization by the LLM tokenizer, which converts characters into a sequence of tokens $\bm{t}_{1:N_{t}}$ for LLM to process. The LLM subsequently generates a compact representation of the text, effectively encoding the tokens into a compressed binary message $\bm{m}\in\{0,1\}^{K}$. Specially, considering a dictionary $\mathcal{D}$ of $\tau$ tokens, the input sequence $\bm{s}$ is first parsed into a sequence of $N_{t}$ tokens $t_{1},t_{2},...,t_{N_{t}}$. Given the first $k$ tokens $\bm{t}_{1:k}$, the $(k+1)$-th token $t_{k+1}$ can be inferred as a predicted probability distribution $\tilde{\rho}(t_{k+1}|\bm{t}_{1:k})$. Here, $\tilde{\rho}(t_{k+1}|\bm{t}_{1:k})$ indicates the LLM estimation of the true distribution $\rho(t_{k+1}|\bm{t}_{1:k})$. The incremental decoding nature in LLM implies that it can accurately predict the probability distribution of the next token based on known ones, thus being capable to give a sub-optimal estimation of true distribution [25]. As shown in Fig. 2, selecting the next character is actually narrowing down the probabilistic interval where the sequence locates, which means the code $\bm{m}$ is determined once the interval is fixed. Starting with $\mathbb{I}_{0}=[0,1)$, the previous interval determined by $\bm{t}_{1:k}$ in step $k$ is defined as $\mathbb{I}_{k}=[l_{k},u_{k})$. Therefore, denoting $p(t_{k+1}=D_{j})=\tilde{\rho}(t_{k+1}=D_{j}|\bm{t}_{1:k})$,

### v2

@@ Abstract
Along with the proliferating research interest in Semantic Communication (SemCom), Joint Source Channel Coding (JSCC) has dominated the attention due to the widely assumed existence in efficiently delivering information semantics. Nevertheless, this paper challenges the conventional JSCC paradigm and advocates for adopting of Separate Source Channel Coding (SSCC) to enjoy the underlying more degree of freedom for optimization. We demonstrate that SSCC, after leveraging the strengths of Large Language Model (LLM) for source coding and Error Correction Code Transformer (ECCT) complemented for channel decoding, offers superior performance over JSCC. Our proposed framework, effectively highlights the compatibility challenges between SemCom approaches and digital communication systems, particularly concerning the resource costs associated with the transmission of high precision floating point numbers. Through comprehensive evaluations, we establish that assisted by LLM-based compression and ECCT-enhanced error correction, SSCC remains a viable and effective solution for modern communication systems. In other words, separate source channel coding is still what we need!
@@
Figure 1: Framework of LLM-based and ECCT-complemented SSCC system.
@@ I Introduction
Semantic communication (SemCom) has garnered significant attention in recent years, with researchers exploring innovative approaches to enhance the efficiency and reliability of information transmission [1]. Generally, SemCom leverages deep learning-based Joint Source-Channel Coding (JSCC) methods to preserve global semantic information and local texture during the transmission process. DeepJSCC [2] pioneers these works by implementing JSCC with feedback and allowing for real-time adaptation to channel conditions. Along with its steady progress, JSCC has been substantially studied, mostly with the optimization objective shifting from bit error rates to the semantic relevance of the transmitted information in SemCom [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. However, albeit the awfully exploded research interest, one critical question remains unsolved: why does the joint approach stand out, as separate source channel coding (SSCC) shall promise a greater degree of freedom from an optimization perspective?
@@
As the terminology implies, SSCC encompasses two decoupled ingredients: source coding and channel coding. The former part lies in effectively compressing the context, and the effectiveness of underlying Deep Neural Networks (DNN)-based predictors, such as Recurrent Neural Networks (RNN)-based DeepZip [15], Long Short Term Memory (LSTM)-based [16, 17] and hybrid DNN-based Dzip [18], have been validated widely in achieving satisfactory text compression. More prominently, Transformer-based [19] and Large Language Model (LLM)-based compression springs up recently [20, 21, 22, 23, 24]. The latest research [25] unveils the equivalence between compression and prediction. In other words, in the general framework where statistical models predict symbols and encoders use predictive probabilities to perform compression, better predictive models lead directly to better compressors [25]. Hence, the astonishing capability of LLM implies the potential for an unprecedented source codec. On the other hand, Error Correction Code (ECC) plays an indispensable role in channel coding. Although some advanced algebraic block codes like Bose–Chaudhuri–Hocquenghem (BCH) codes [26], Low-Density Parity-Check (LDPC) codes [27] and Polar codes [28], can somewhat ensure the reliability of transmission, the efficient decoding of ECC is an unresolved difficulty. Recently, DNNs have started to demonstrate their contribution in channel coding. For example, deep learning models are implemented to achieve Belief Propagation (BP) decoding [29, 30, 31], while a model-free Error Correction Code Transformer (ECCT) for algebraic block codes [32] contributes to the enhancement of decoding reliability.
@@
In this paper, on top of an LLM-based Arithmetic Coding (LLM-AC) system, the proposed SSCC framework integrates fine-grained, semantics-aware probability modeling and encoding with ECCT-enhanced channel decoding, thus forming a closed-loop optimization framework. To the best of our knowledge, this work represents the very first comprehensive integration of LLM-based compression and ECCT-complemented channel decoding for a holistic SemCom architecture. Through extensively showcasing the performance superiority over JSCC, we argue this performance improvement primarily arises after tackling the underlying incompatibility between conventional SemCom approaches [3, 7, 8, 9, 10, 11, 12, 13, 14] and digital communication architectures [33]. Particularly, those approaches simply assume the deliverability of encoded semantic feature vectors while neglecting the energy costs associated with transmitting high-precision floating point numbers [33]. However, further quantization [5, 6] and digital modulation can compromise the widely assumed existence of performance superiority in JSCC. Meanwhile, in contrast to the direct utilization of the astonishing semantic interpretation capability [34, 35, 36], the deployment of LLMs focuses on the compression and encoding of text to squeeze the largely untapped redundancy. Therefore, our work is also significantly different from existing integrations of Generative AI (GAI) and SemCom [37, 38, 39, 40, 41, 42, 43]. Furthermore, the adoption of ECCT boosts the effectiveness of SSCC in specific cases. In a nutshell, our comprehensive evaluation of LLM and ECCT-based SSCC shows that separate source channel coding is still what we need!
@@
The rest of this paper is organized as follows. Section II introduces the SSCC system model, while its key components are enumerated in Section III. Numerical results are presented in Section IV to show the performance superiority of the proposed SSCC system. Finally, Section V concludes this paper with discussions on future works. For convenience, we list the major notations of this paper in Table I.
@@
Notation Definition
$\bm{s},\bm{\hat{s}}$ Source text sequence and its reconstructed version at the receiver side
$\bm{t},\bm{\hat{t}}$ Token sequence and its reconstructed version at the receiver side
@@
$\mathcal{D},D_{i},\tau$ Source code dictionary, its $i$-th character and total vocabulary size
$\mathbb{I}_{k},l_{k},u_{k}$ Probability interval at step $k$ of source coding and its corresponding lower bound and upper bound
$\bm{m},\bm{\hat{m}}$ Source-coded message vector and its channel-decoded version
$\lambda$ Probability interval represented in decimal form
$N,K$ Codeword length and message length of error correction code $C_{e}(N,K)$
$\bm{G,H}$ Generator matrix and parity check matrix
$\bm{x},\bm{x}_{b},\bm{x}_{s}$ Transmitted codeword vector encoded by the channel coder, its binary representation (taking values in 0,1)and its signed representation (taking values in -1,+1)
$\bm{\hat{x}},\bm{\hat{x}}_{b}$ Soft approximation of codeword, and its binary representation
$\mathcal{N}(\cdot,\cdot),\sigma_{n}$ Gaussian distribution and noise standard deviation
@@
$\bm{z},\bm{\tilde{z}},\bm{\hat{z}}$ Additive Gaussian noise, corresponding multiplicative noise and ECCT prediction result
$\bm{y},\bm{y}_{b},\bm{{\tilde{y}}}$ Received noisy codeword, its binary representation and pre-processed version
$\texttt{syn}(\cdot)$ Syndrome computation function defined in ECCT
$f(\cdot)$ ECCT decoding function
$\bm{W}$ Learnable embedding matrix for high-dimensional mapping
$\bm{g}(\cdot)$ Code-aware self-attention mask function
@@ II System Model
  • Source Encoding: The input text sequence denoted as $\bm{s}_{1:N_{s}}$ undergoes a source encoder that converts characters into a compressed binary message $\bm{m}\in\{0,1\}^{K}$. During source encoding, Arithmetic Coding (AC) can be leveraged for effective compression here. For LLM-based processing, an intermediate result (i.e., a sequence of tokens $\bm{t}_{1:N_{t}}$) can be obtained during the transformation from $\bm{s}$ to $\bm{m}$.
  • Channel Encoding and Modulation: The message $\bm{m}$ is then encoded via a LDPC code $C_{e}(N,K)$, which is selected for its excellent error-correction capabilities and compatibility with iterative decoding algorithms, as mentioned in [32]. The encoding process employs a generator matrix $\bm{G}$ to transform the message in $\bm{m}$ to a codeword $\bm{x}\in\{0,1\}^{N}$. The parity check matrix $\bm{H}$, which satisfies $\bm{G}\cdot\bm{H}^{T}=0$ and $\bm{H}\cdot\bm{x}=0$ is a key component of the LDPC decoding process. Afterwards, Binary Phase Shift Keying (BPSK) modulation maps the binary codeword $\bm{x}$ to a sequence of symbols $\bm{x}_{s}\in\{0,1\}^{N}$ suitable for transmission over the wireless channel. Notably, other error correction codes such as Polar codes [28] can be applied as well.
@@
  • Demodulation and Channel Decoding: BPSK demodulation recovers a binary codeword $\bm{\hat{x}}_{b}\in\{0,1\}^{N}$ from $\bm{\hat{x}}$. Subsequently, channel decoder reconstructs the message $\bm{\hat{m}}\in\{0,1\}^{K}$ from $\bm{\hat{x}}_{b}$. In contrast to conventional approaches that employ either hard-decision (e.g., bit-flipping algorithm) or soft-decision (e.g., sum-product algorithm) algorithms to decode LDPC codewords transmitted through the channel, some complementary decoding modules, such as ECCT, can be applied prior to demodulation to enhance the decoding performance. Notably, ECCT can provide an estimation of the transmitted codeword $\bm{\hat{x}}_{b}$, denoted as $\bm{\hat{x}}$, while subsequent demodulation and information bits extraction are then performed on the estimated codeword $\bm{\hat{x}}$.
  • Source Decoding: The recovered message $\bm{\hat{m}}$ is ultimately decoded by the source decoder, which reconstructs the original text sequence $\bm{s}$ from the message, effectively reversing the encoding process. Similar to the encoder, the decoder can implement arithmetic decoding.
In comparison, JSCC often leverages an end-to-end DNN to implement source and channel codecs. Here, the terminology “end-to-end” implies the generally joint training of source and channel codes, as adopted in most works. The further details on JSCC can be found in [1] and the references therein. In the following section, we will address how to leverage the strength of LLM to enhance text compression and reconstruction, combined with the robustness of ECCT-complemented LDPC codes for error correction, as shown in Fig. 1.
@@ III-A LLM-based Source Coding
Given a source distribution $\rho$, lossless compression aims to encode a text sequence $\bm{s}$ sampled from $\rho$ into a binary code $\bm{m}=C_{s}(\bm{s})$ of minimal possible length with no loss of original information. According to Shannon’s source coding theorem [44], the optimal expected bit length is $L_{\min}=\mathbb{E}_{\bm{s}\sim\rho}[-\log_{2}\rho(\bm{s})]$. To obtain such optimal length, arithmetic coding [45, 46], a form of entropy encoding, is typically adopted contingent on a probabilistic model over $\rho$ or its marginal distribution. Arithmetic coding implies that frequently used characters are stored with fewer bits while rarely occurred characters correspond to more bits, resulting in fewer bits used in total.
In particular, the input text sequence denoted as $\bm{s}$ undergoes tokenization by the LLM tokenizer, which converts characters into a sequence of tokens $\bm{t}$ for LLM to process. The LLM subsequently generates a compact representation of the text, effectively encoding the tokens into a compressed binary message $\bm{m}\in\{0,1\}^{K}$. Specially, considering a dictionary $\mathcal{D}$ of $\tau$ tokens, the input sequence $\bm{s}$ is first parsed into token sequence $\bm{t}$. Given the first $k$ tokens $\bm{t}_{1:k}$, the $(k+1)$-th token $t_{k+1}$ can be inferred as a predicted probability distribution $\tilde{\rho}(t_{k+1}|\bm{t}_{1:k})$. Here, $\tilde{\rho}(t_{k+1}|\bm{t}_{1:k})$ indicates the LLM’s estimation of the true distribution $\rho(t_{k+1}|\bm{t}_{1:k})$. The incremental decoding nature in LLM implies that it can accurately predict the probability distribution of the next token based on known ones, thereby providing a sub-optimal estimation of the true distribution [25]. As shown in Fig. 2, selecting the next character is actually narrowing down the probabilistic interval where the sequence is located, which means the code $\bm{m}$ is determined once the interval is fixed. Starting with $\mathbb{I}_{0}=[0,1)$, the previous interval determined by $\bm{t}_{1:k}$ in step $k$ is defined as $\mathbb{I}_{k}=[l_{k},u_{k})$. Therefore, denoting $p(t_{k+1}=D_{j})=\tilde{\rho}(t_{k+1}=D_{j}|\bm{t}_{1:k})$,